OpenStudy (anonymous):

what is the derivative of: 2 square x (not sure how to do) and (1-2x)/x^3 I get 3x^3(-1+2x) not sure if correct

OpenStudy (anonymous):

sorry for the last one my answer is 3x^2(-1+2x)

OpenStudy (anonymous):

u can rewrite square of x by x^1/2

OpenStudy (anonymous):

2(sq.x)=2(x^1/2)

OpenStudy (anonymous):

multiply the exponent by the degree and subtract 1 from the exponent

OpenStudy (anonymous):

\[2\sqrt{?}\] sorry I can't get it in corrrectly the ? should be the x

OpenStudy (amistre64):

take the 2 out and derive sqrt(x); or x^(1/2) 1/2sqrt(x) now put the 2 back in: 2/2sqrt(x) = 1/sqrt(x)

OpenStudy (amistre64):

(1-2x)/x^3 use quotient rule :) x^3(-2) - (1-2x)(3x^2) -------------------- x^6

OpenStudy (anonymous):

\[\frac{d}{dx}[ \frac{(1-2x)}{x^3}] \] \[= \frac{d}{dx}[ (1-2x)x^{-3}]\] \[= (1-2x)\frac{d}{dx}[ x^{-3}] + x^{-3}\frac{d}{dx}[ (1-2x)]\] \[ = (1-2x)(-3x^{-4}) + x^{-3}(-2)\] \[ = -3x^{-4} + 6x^{-3} - 2x^{-3}\] \[ = \frac{4x-3}{x^4} \]

OpenStudy (amistre64):

-2x^3 -3x^2+6x^3 4x^3 - 3x^2 (4x-3)x^2 -------- x^4 x^2 4x-3 ----- maybe?? x^4

OpenStudy (anonymous):

thanks, i have to go for now , i will check back in about and hour

OpenStudy (amistre64):

polpak did it good :) or we messed up toghter lol

OpenStudy (anonymous):

I prefer to ignore the quotient rule and use product rule with a negative exponent.

OpenStudy (amistre64):

potato<>potato :)

OpenStudy (anonymous):

I thought potato = potato ;p

OpenStudy (radar):

There is another way to look at mom. (1-2x)/x^3 can be rewritten to appear as (1-2x)(x^-3) \[(1-2x) *x ^{-3}=x ^{-3}-2x ^{-2}\]

OpenStudy (radar):

This may be in a more familiar form you can now simply take the derivative to be: \[-3x ^{-4}+4x ^{-3}\] simplifying will obtain the results of polpak and amistre64